OpenStudy (anonymous):

Is the series convergent or divergent? I don't know where to start. I'll type it below.

OpenStudy (anonymous):

You need to perform a convergence test on the given series.

OpenStudy (anonymous):

You didn't write out the series, but I will say that there are many different convergence tests.

OpenStudy (anonymous):

Here are a few: Ratio test Integral test Comparison test etc. http://en.wikipedia.org/wiki/Convergence_tests

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}e ^{n}/n ^{2}\]

OpenStudy (anonymous):

Do divergent test, does it go to zero as n approach infinity?

OpenStudy (anonymous):

This is a problem in the chapter before all of the tests you listed.

OpenStudy (anonymous):

Well its fairly obvious that this series diverges.

OpenStudy (anonymous):

as n approaches infinity, this ratio does not tend to zero, hence it is diverges

OpenStudy (anonymous):

Since the exponential function grows much faster than n^2

OpenStudy (anonymous):

is there another way to show this or I can only look at the graphs?

OpenStudy (anonymous):

Divergence test is usually first test performed to see whether a series converges or diverges

OpenStudy (anonymous):

in this case divergent test shows infinity/infinity, so I'm not so sure.

OpenStudy (zarkon):

Use L'Hospitals rule or the taylor expansion of e^n to show divergence as n->infinity

OpenStudy (anonymous):

Do L'Hopital you will see e^x grows much faster than x^2

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

Ratio test: \[\large L = \lim_{n\to \infty}\frac{\frac{e^{n+1}}{(n+1)^2}}{\frac{e^n}{n^2}} = e\] L > 1 Series diverges